HUBBLE SPACE TELESCOPE DAILY REPORT #870 PERIOD COVERED: 0800 EST Daily Status Report as of 111/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WFC 4105 (HST Medium Deep Survey - Cycle 1 High Latitude)) The Wide Field Camera was used to make one image for a high latitude medium deep sky survey. This image was taken in parallel with the prime science observations. The image was received and displayed in the OSS. It was reported to look okay, however, it did have a satellite trail across the image. 1.2 Completed PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make one image each for internal flat calibration images. These are looking for new forms of contamination (i.e., daises and measles), and the return of QEH (quantum efficiency hysteresis). Both images were received and displayed in the OSS. They both were reported to look okay. 1.3 Completed PC 4778 (Recent Mass Ejection From Planetary Nebulae Nuclei) The Planetary Camera was used to make images of the central star in the protoplanetary nebula EGB6 (i.e., star 0950+139). The star has been discovered to have a companion at a separation of 0.18 arcseconds. Three different recentering events were reported to have occurred during the observations due to day/night transitions. The observations were received and displayed for the principal investigator. They were reported to look okay. 1.4 Completed PC 4140 (WFPC Cycle 2 Calibration: Rapid Internal Monitor - Part 2)) The Planetary Camera was used to make three internal flat calibration observations using filter F439W to check for new forms of contamination. All three observations were received and displayed in the OSS. They were reported to look okay. 1.5 Completed FOS/RD 3660 (The Level of Ionization and Chemical Composition of QSOs) The Faint Object Spectrograph was used to make two observations of the broad absorption line regions within quasar Q0932+501 to help confirm its chemical composition. The onboard binary target acquisition succeeded, with eight groups in the final file. Two separate recentering events were reported to have occurred during the observations. The last half of the observations occurred during LOS, and the science data still remains on the tape recorder. No problems were reported. 1.6 Completed WFC 3917 (HST Medium Deep Survey: Cycle Two Part 1) Both the FOC (f/96) and the Wide Field Camera were used in parallel to do a high latitude sky survey. Three images were taken with the WFC in parallel with the other proposal. Two different recentering events occurred during the observations. The first observations was received and displayed. The other two observations occurred during LOS, and will be analyzed later. 1.7 Completed HSP/UV2 4801 (Quasi-Periodic Oscillations in AM Hercules Binaries)) The High Speed Photometer (Ultraviolet side) was used to study high speed oscillations in the variable star VV Puppis (Which is a AM Hercluis class variable star). The observations were executed as scheduled, and no problems were reported. 1.8 Completed FOS/BL 4054 (Probing the Nuclear Regions of the Seyfert Galaxy NGC-5548) The Faint Object Spectrograph was used to make spectral observations of the nuclear regions in Seyfert galaxy NGC-5548 to study the time variability of its continuum fluxes, and emission lines of Helium II and Carbon IV. The observations were executed as scheduled, and no problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Note: The MOSES/NASA power engineers have received the 116 SMS and found that no SA slew are necessary, given the small changes recommended by the STScI. The EPS will be power positive throughout the SMS. Further, the STScI will , starting with the 123 SMS, recommend SA slews only when a minimum of 5.8 degree change is necessary. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4105 (HST Medium Deep Survey: Cycle 1 High Latitude) 3.2 Complete Two Sets of HSP/UV2 4801 (Quasi-Periodic Oscillations in AM Herculis Binaries) 3.3 Complete Two Sets of FOC/96 4800 (The Stellar Content of Wolf-Rayet Galaxies (FOC images)) 3.4 Complete Two Sets of PC 4106 (HST Medium Deep Survey: Cycle 1 Low Latitude) 3.5 Complete FOS/RD 4220 (Imaging and Spectroscopy of Seyfert nuclei) 3.6 Complete WFC 1297 (GTO High Latitude Parallel Program) 3.7 Complete FOS/BL 4054 (Probing the Nuclear Regions of the Seyfert Galaxy NGC-5548) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-